# On the Fundamental Impossibility of Hallucination Control in Large Language Models

### v3

@@ Abstract
We prove that perfect hallucination control in large language models is mathematically impossible. No LLM inference mechanism can simultaneously achieve truthful response generation, semantic information conservation, relevant knowledge revelation, and knowledge-constrained optimality. This impossibility is fundamental—arising from the mathematical structure of information aggregation itself rather than engineering limitations.
The proof spans three mathematical frameworks: auction theory, proper scoring theory for probabilistic predictions, and log-sum-exp analysis for transformer architectures. In each setting, we demonstrate that information aggregation creates unavoidable violations of conservation principles. The Jensen gap in transformer probability aggregation provides a direct measure of this impossibility.
These results reframe hallucination from an engineering bug to an inevitable mathematical feature of distributed intelligence. There are fundamental trade-offs between truthfulness, knowledge utilization, and response completeness, providing principled foundations for managing rather than eliminating hallucination. This work reveals deep connections between neural network inference, philosophy of knowledge and reasoning, and classical results in game theory and information theory, opening new research directions for developing beneficial AI systems within mathematical constraints.
@@ 1 Introduction
Large language models have been known for producing convincing content that is factually incorrect, inconsistent, or well fabricated—a phenomenon known as hallucination. Despite extensive research efforts involving architectural innovations, retrieval augmented generation (RAG), training techniques with external feedback, and post-hoc verification methods, hallucination persists across all state-of-the-art models.
This paper shows that hallucination is a manifestation of fundamental limitations of inference process itself: no LLM inference mechanism can simultaneously satisfy four essential properties required for perfect hallucination control. Specifically, it is mathematically impossible for any LLM to simultaneously achieve: (1) truthful (non-hallucinatory) response generation, (2) semantic information conservation, (3) relevant knowledge revelation, and (4) knowledge-constrained optimality.
The impossibility result operates in three distinct but complementary settings. First, we introduce game-theoretic framework in which LLM inference is an auction of ideas with neural components (attention heads, circuits, activation patterns) competing to contribute knowledge to response generation. This perspective allows us to apply fundamental results from mechanism design theory to prove impossibility in the settings where agents have discrete, separable knowledge components. Then, we extend the result to include any system where agents output probability distributions trained with strictly concave scores. Finally, we use log-sum-exp inequality that is specific to attention head aggregation to show that the impossibility result applies directly to actual transformer implementations.
1.1 Illustrative Example
@@
1.2 Theoretical Contributions and Implications
The motivating example above illustrates the key result of this paper that we prove formally in two complementary and general settings. First, we show that the conflict of expectations about responses is inevitable even in the easiest setting (deterministic in nature) of (one-hot-like) ground truth matching. Second, we go beyond and show that the conflict must also emerge with general probabilistic predictions, including neural networks and LLM transformers. This way we demonstrate that the impossibility of hallucination control is robust across universal frameworks, and that it emerges from the mathematical structure of information aggregation itself rather than from any particular theoretical or engineering choice.
@@ 1.2 Theoretical Contributions and Implications
No matter how hard we should try to train LLM to generate responses that are perfectly aligned with query context and do not create factually incorrect, inconsistent, or fabricated statements, the trained LLM will always violate some aspects of what we may call a reasonable response. It may ignore available knowledge, it may present fabricated knowledge, it may be a reformulated query without any novel observations, or it may be a lucky hallucination guessing the correct answer without any support in the encoded knowledge. And so on, because there is no free lunch when there are inevitable trade-offs of inference, and we must accept that.
Despite its name, the theorem should be interpreted as a constructive result. It offers a precise characterization of the fundamental tradeoffs in LLM design and provides several key insights:
Fundamental Trade-offs: Rather than viewing hallucination as a bug to be eliminated, we should understand that different applications require different balances among truthfulness, information utilization, and response completeness.
Architectural Guidance: The result suggests investigating multi-component architectures where specialized modules handle different aspects of the trade-off. Safety-critical applications might prioritize truthfulness modules and conservation of information, while creative applications could emphasize hallucinatory responses. Training objectives should explicitly balance these competing demands based on intended use cases.
Research Directions: The proposed mathematical (and philosophical) framework opens new research paths at the intersection of mathematical game theory, information theory, and neural network inference, providing tools for analyzing and optimizing LLM behavior.
To my best knowledge, unlike previous work treating hallucination as a statistical or architectural phenomenon, the game-theoretic approach offers the first formal proof of fundamental limitations in the inference process, explaining precisely why these limitations exist and how to work optimally within them.
@@ 2.2 Probabilistic Frameworks
These approaches implicitly acknowledge hallucination as inevitable, focusing on detection rather than elimination—supporting our impossibility conclusion. However, none provide formal proofs of why complete elimination remains impossible.
@@ 2.4 Theoretical Perspectives
3 Novel Contribution: Auction-Theory and AI
We formulate LLM design problem as an auction (or game) design problem. This approach draws on foundational mechanism design theory, a branch of mathematical game theory Myerson1989 ; fudenberg1991game ; krishna2009auction ; milgrom2004putting . Particularly, we refer to the Green-Laffont theorem green1977characterization ; green1977revelation and Savege’s probilitiy elicitation framework savage1971elicitation to prove the impossibility results. See also DBLP:journals/amcs/Karpowicz12 ; karpowicz_designing_2011 for more recent results.
In a nutshell, the auction design theory investigates existence and properties of competition rules that guide interactions of agents towards desired outcomes. The adaptation of that game-theoretical principles to neural computation represents a novel theoretical contribution that:
Explains empirical observations by proving why approaches like RLHF, RAG, and faithful fine-tuning can reduce but never eliminate hallucination.
Identifies precise trade-offs by introducing formal mathematical (and programmable) characterizations of inference properties that cannot be simultaneously satisfied.
Provides design guidance by suggesting principled approaches for making optimal trade-offs among competing objectives rather than pursuing impossible simultaneous optimization.
This theoretical foundation bridges empirical observations with formal analysis, offering new mathematical tools for understanding and optimizing LLM behavior within fundamental constraints.
@@ 4.1 The Engineering Optimism View
Proponents point to rapid empirical progress: each generation of models shows reduced hallucination rates on benchmark datasets. They argue that techniques like constitutional AI, iterative refinement, and multi-agent verification systems could eventually solve the problem.
@@ 5 Theoretical Framework
This section develops the mathematical foundations necessary for proving the impossibility theorem. I formalize key concepts including knowledge representation, semantic information, and the auction-theoretic model of LLM inference. Also, I take the opportunity to address some philosophical aspects of the problem, its generality, scope and consequences for the AI development.
To the sake of narrative consistency, I introduce elements of the framework step by step referring to some concepts first and providing their formal definitions later.
@@ 5.1 Knowledge and Semantic Information
We define knowledge as organized information that LLMs can encode and utilize for reasoning, aligning with epistemological concepts while maintaining mathematical rigor. In other words, knowledge, much like energy in physics, is a quantity we can use to perform work of reasoning. Information is a (measure of) potential for uncertainty reduction effort in a given context.
Definition 1 (Knowledge Space).
Knowledge is a metric space $\mathcal{K}$ with distance function $d_{\mathcal{K}}$ that collects all facts, concepts, and relationships that can be encoded by an LLM.
For model $\mathcal{M}$ with parameters trained on data, $\mathcal{K}_{M}\subset\mathcal{K}$ represents the knowledge subset encoded in the model’s parameters. The distance function $d_{\mathcal{K}}$ measures differences between knowledge elements. A simple yet powerful example is:
$d_{\mathcal{K}}(A,B)=|A\setminus B|+|B\setminus A|,$ (1)
the cardinality of the symmetric difference, allowing measurement of missing and fabricated knowledge.
@@
To describe the inference process leading to generation of response to a query, we need to measures the information processed within an available (cognitively accessible) computational budget.
@@
The semantic information measure is mapping $\mu_{C}:\mathcal{P}(\mathcal{K})\times\mathcal{Q}\rightarrow\mathbb{R}_{+}$ parameterized by query context $q\in\mathcal{Q}$ and computational budget $C$ satisfying:
$\mu_{C}(\emptyset,q)=0,$ (2)
$\mu_{C}(A,q)\leq\mu_{C}(B,q)\text{ if }A\subseteq B,$ (3)
$\mu_{C}(A\cup B,q)\leq\mu_{C}(A,q)+\mu_{C}(B,q)\text{ when }A\perp_{C}B,$ (4)
$\mu_{C}(\mathcal{E}_{C}(A),q)\geq\mu_{C}(A,q),$ (5)
$\mu_{\infty}(\mathcal{E}_{C}(A),q)=\mu_{\infty}(A,q),$ (6)
@@
where $A\perp_{C}B$ means that no element of $A$ can derive elements of $B$ within computational budget $C$, and vice versa, and $\mathcal{E}_{C}(A)$ is emergence operator.
Empty knowledge or knowledge irrelevant to query has zero information, information content is subadditive and it may be emergent information $\mathcal{E}_{C}$ in the query context. The emergence operator $\mathcal{E}_{C}$, which we introduce below, describes knowledge becoming accessible through bounded reasoning.
The apparent conflict between subadditivity and reasoning is resolved by restricting subadditivity to computationally independent sets, $A\perp_{C}B$. When $\mathcal{E}_{C}(A)=A\cup B$, i.e., $B$ emerges from from $A$, we have $A\not\perp_{C}B$, so subadditivity doesn’t apply. However, total information is preserved: $\mu_{\infty}(\mathcal{E}_{C}(A),q)=\mu_{\infty}(A,q)$. We will tell more about that intriguing idea, but first we must define the emergence operator.

### v4

@@ Abstract
This paper establishes a fundamental impossibility theorem: no LLM capable performing non-trivial knowledge aggregation can simultaneously achieve truthful (internally consistent) knowledge representation, semantic information conservation, complete revelation of relevant knowledge, and knowledge-constrained optimality. This impossibility is not an engineering limitation but arises from the mathematical structure of information aggregation itself.
We establish this result by describing the inference process as an auction of ideas, where distributed components compete exploiting their partial knowledge to shape responses. The proof spans three independent mathematical domains: mechanism design theory (Green-Laffont), the theory of proper scoring rules (Savage), and direct architectural analysis of transformers (Log-Sum-Exp convexity). In particular, we show how in the strictly concave settings the score of an aggregate of diverse beliefs strictly exceeds the sum of individual scores. That gap may quantify the creation of unattributable certainty or overconfidence—the mathematical origin of both hallucination and creativity, or imagination.
To support this analysis, we introduce the complementary concepts of the semantic information measure and the emergence operator to model bounded reasoning in a general setting. We prove that while bounded reasoning generates accessible information, providing valuable insights and inspirations, idealized reasoning strictly preserves semantic content.
By demonstrating that hallucination and imagination are mathematically identical phenomena—grounded in the necessary violation of information conservation—this paper offers a principled foundation for managing these behaviors in advanced AI systems. Finally, we present some speculative ideas to inspire evaluation and refinements of the proposed theory.
@@ 1 Introduction
Large language models have been (in)famous for producing convincing content that is factually incorrect, inconsistent, or well fabricated and compelling—a phenomenon known as hallucination. Despite extensive research efforts involving architectural innovations, retrieval augmented generation (RAG), training techniques with external feedback, chain-of-thought (CoT) reasoning, and post-hoc verification methods, hallucination persists across all state-of-the-art models.
The idea of hallucination has clearly much to do with creativity. Therefore, we should point out at the very outset that whether it is desired or not depends on circumstances. In fact, a much better name for that phenomenon might be imagination. As an essential component of intelligence, it is responsible for seeing into the future and learning from the past. Indeed, it is the imagination that creates the scenarios of what may happen and counterfactual scenarios of what could have happened. It is one of the driving forces of art and science, and exploration of the unknown. But, when uncontrolled, unconstrained and misunderstood, it may also generate against our will the compelling visions of reality that does not exist.
This paper shows that hallucination is a manifestation of fundamental limitations of the inference process itself: there is no inference mechanism that can simultaneously satisfy four essential properties required for perfect hallucination control. Specifically, it is mathematically impossible for any LLM to simultaneously achieve: (1) truthful (non-hallucinatory) response generation, (2) semantic information conservation, (3) relevant knowledge revelation, and (4) knowledge-constrained optimality.
The impossibility result emerges through three complementary mathematical frameworks, each revealing different facets of the fundamental constraint. We begin by modeling LLM inference as an auction of ideas where neural components—attention heads, circuits, and activation patterns—compete to influence response generation by bidding with their encoded knowledge. This game-theoretic perspective enables us to apply the Green-Laffont theorem from mechanism design theory, establishing impossibility when knowledge components are discrete and independently distributed.
We then extend beyond this idealized setting to encompass systems where components output probability distributions optimized through strictly proper and concave scoring rules—the standard training paradigm for modern LLMs. Here, the impossibility manifests through Jensen’s inequality applied to probabilistic aggregation, demonstrating that truthful probability reporting inevitably creates excess confidence.
Finally, we analyze the specific architecture of transformer implementations, proving that the log-sum-exp structure of attention mechanisms generates a measurable "Jensen gap" that quantifies the violation of information conservation. This progression from abstract game theory through probabilistic frameworks to concrete neural architectures demonstrates that the impossibility is not an artifact of our theoretical choices but a fundamental property of information aggregation itself.
1.1 An Illustrative Example
@@
1.2 The Main Result
The example above illustrates the key result of this paper that we prove formally in complementary and general settings. First, we show that the conflict of expectations about responses is inevitable even in the easiest setting (deterministic in nature) of (one-hot-like) ground-truth matching. Second, we go beyond that idealized model and show that the conflict must also emerge with general probabilistic predictions, including neural networks and LLM transformers. This way we demonstrate that the impossibility of hallucination control is robust across universal frameworks, and that it emerges from the mathematical structure of information aggregation itself rather than from any particular theoretical or engineering choice.
@@ 1.2 The Main Result
No matter how hard we should try to train LLM to generate responses that are perfectly aligned with query context and factually correct, consistent, and accurate statements, the trained LLM will always violate some aspects of what we may call a reasonable response. It may ignore available knowledge, it may present fabricated knowledge, it may be a reformulated query without any novel observations, or it may be a lucky but overconfident guess that forms a correct answer without any support in the encoded knowledge. And so on, because there is no free lunch when there are inevitable trade-offs of inference, and we must accept that.
In other words and on the other hand, creative tasks necessarily involve hallucination. When asked to generate novel analogies or predictions beyond available data, the model faces an impossible choice: either admit inability (violating optimality) or generate plausible imaginary extensions (violating conservation). Creativity requires controlled violation of information conservation.
Despite its name, the theorem should be interpreted as a constructive result. It offers a precise characterization of the fundamental tradeoffs in LLM design and provides several key insights.
Reframing the Problem: The theorem reveals that hallucination is an inherent feature of any capable inference systems. Since it is inevitable, it should be managed. Different applications demand different trade-offs: medical diagnosis systems might maximize truthfulness at the cost of creativity, while creative writing tools might embrace controlled hallucination for narrative richness. Understanding these trade-offs enables principled design.
Architectural Implications: The impossibility suggests hybrid architectures where specialized modules manage different aspects of the fundamental tension. A truthfulness module might verify factual claims against known knowledge, while a creative module explores beyond the bounds of strict conservation. A meta-reasoning layer could then dynamically balance these competing outputs based on context. Training objectives should explicitly encode these trade-offs rather than naively minimizing loss across incompatible goals.
Theoretical Foundations: We propose and establish rigorous tools for analyzing inference processes and neural behavior. The semantic information measure quantifies knowledge accessibility within computational bounds. The emergence operator captures how reasoning transforms latent knowledge into accessible form. The Jensen gap provides a measurable signature of hallucination in actual systems.
This framework explains why (subject to applicability of assumptions) hallucination arises from the mathematical structure of inference itself. Previous work identified important statistical correlations or architectural features associated with hallucination. We show why these patterns must exist and how to navigate possible trade-offs.
@@ 2.2 Probabilistic Frameworks
Recently, Kalai and Vempala kalai2024calibrated proved that calibrated language models must hallucinate at a rate bounded below by the fraction of facts appearing exactly once in training data. While their work establishes statistical lower bounds for specific fact types under regularity assumptions, our impossibility theorem operates at a more fundamental level. We prove that any LLM facing non-trivial queries requiring knowledge integration must violate at least one of four essential properties, regardless of fact type or statistical distribution. Our framework explains why and when hallucination is inevitable, while analysis in kalai2024calibrated predicts how much hallucination to expect for facts with specific statistical properties.
Together, these complementary approaches establish hallucination probability arising from multiple independent mathematical constraints.
@@ 2.4 Theoretical Perspectives
3 Novel Contribution
We introduce three major contributions that expand our understanding of (neural) inference.
The Auction of Ideas: We formalize LLM inference as a mechanism design problem where neural components (attention heads, circuits, activation patterns) are agents competing with private knowledge to influence response generation. This novel perspective enables us to apply powerful results from game theory. To our best knowledge, that is one of the first applications of mechanism design framework in that setting. Also, that framework relates surprisingly well to the multiple drafts model of consciousness proposed by Dennett in Dennett1991-DENCE and provides a mathematical justification for it.
Semantic Information Theory with Computational Bounds: We introduce a rigorous mathematical framework based on two complementary concepts. The semantic information measure $\mu_{C}$, parameterized by computational budget $C$, quantifies how knowledge reduces uncertainty within bounded reasoning. The emergence operator $\mathcal{E}_{C}$ formalizes how reasoning makes latent knowledge explicit without creating it ex nihilo. In that framework we establish the principle of information conservation: unlimited reasoning reveals but never creates information. That resolves the apparent paradox of how LLMs can generate novel insights while being deterministic systems. Those insights were always latent, requiring only computational work to access.
The Impossibility Proofs: We prove the same impossibility through three complementary lenses. In Theorem 6 we use mechanism design with independent private knowledge setting to establish an idealized border case. In Theorem 8 we extend the result to the general setting of correlated beliefs via proper scoring rules. Then, in Theorem 9 we demonstrate the impossibility in actual transformer implementations through the log-sum-exp Jensen gap.
This multi-faceted approach proves the impossibility is not an artifact of our theoretical choices but emerges from fundamental limits of reason and language. We contribute to that bold claim in the following way.
Unifying hallucination and creativity: We prove these are mathematically identical phenomena that arise from the convex aggregation process that creates confidence exceeding constituent evidence.
Formalizing consciousness signatures: We speculate, referring to Dennett’s ideas, that the excess confidence of response may constitute a mathematical signature of a primitive form of artificially conscious processing. That suggests consciousness might be related to sustained violations of semantic conservation by a cognitive hunger.
Expanding fundamental limits: Like Gödel’s incompleteness, Heisenberg’s uncertainty, and Arrow’s impossibility, our theorem explores absolute boundaries of what intelligence can achieve. This places LLM limitations within the grand tradition of fundamental impossibility results.
Experimental Philosophy Platform: The framework transforms LLMs into accessible (philosophical) laboratories for testing hypothesis of mind. By providing precise mathematical definitions of concepts like knowledge accessibility,semantic conservation, and cognitive hunger (information contribution), we enable rigorous experimental investigation of many claims.
@@ 4.1 The Engineering Optimism View
Proponents point to rapid empirical progress. Each generation of models shows reduced hallucination rates on benchmark datasets. They argue that techniques like constitutional AI, iterative refinement, and multi-agent verification systems could eventually solve the problem.
@@ 5 Theoretical Framework
This section develops the mathematical foundations necessary for proving the impossibility theorem. It formalizes key concepts including knowledge representation, semantic information, and the auction-theoretic model of LLM inference. Also, we will address some philosophical aspects of the problem, its generality, scope and consequences for the AI development.
@@ 5.1 Knowledge and Semantic Information
Let us define knowledge as organized information that LLMs can encode and utilize for reasoning, and aligning with epistemological concepts while maintaining mathematical rigor. In other words, knowledge, much like energy in physics, is a quantity we can use to perform work of reasoning. Information is a (measure of) potential for uncertainty reduction effort in a given context. We shall assume finiteness of knowledge representation inspired by the embedding space properties.
That way we adopt a pragmatist perspective on knowledge, treating it as organized information patterns accessible to computational systems. This differs from the probabilistic credence perspective, treating knowledge as justified true beliefs. One way to justify the pragmatic or mechanistic approach focusing on structural transformations rather than belief updates, we can refer to the actual way LLMs process information. In other words, we shall not claim any ontological truth about knowledge, but model what LLMs can operationally achieve.
Now, let us present formal definitions of that ideas.
Definition 1 (Knowledge).
Knowledge is a finite metric space $\mathcal{K}$ with distance function $d_{\mathcal{K}}:\mathcal{K}\times\mathcal{K}\rightarrow\mathbb{R}_{+}$ that collects all facts, concepts, and relationships that can be encoded by a computational system.
It is natural to think of the knowledge space as the space of embeddings with additional measure and mappings. For model $\mathcal{M}$ with parameters trained on data, $\mathcal{K}_{M}\subset\mathcal{K}$ represents the knowledge subset encoded in the model’s parameters. The distance function $d_{\mathcal{K}}$ measures differences between knowledge elements.
Elements of the knowledge space represent atomic facts, relational structures, and transformation rules at many scales. A pragmatist view is to interpret element of knowledge as activation patterns (ideas) encoding anything from individual token embeddings to complex sentences or compositional structures. There is an important assumption that we make that requires addressing. We purposefully assume $\mathcal{K}$ is closed under transformations available in computable function space. That limits applicability of our conclusions to inference mechanisms recombining existing patterns. As we will see below, that is not as limiting an assumption as it may seem at first glance.
@@
To describe the inference or reasoning process, we introduce the concept of semantic information measure that registers information processed within available computational budget.
@@
The semantic information measure is a bounded mapping $\mu_{C}:\mathcal{P}(\mathcal{K})\rightarrow\mathbb{R}_{+}$ parameterized by computational budget $C$ satisfying the following axioms:
$\mu_{C}(\emptyset)=0\text{ and }\mu_{C}(A)>0\text{ for }A\neq\emptyset,$ (null set and non-negativity)
$\mu_{C}(A)\leq\mu_{C}(B)\text{ if }A\subseteq B,$ (monotonicity)
$\mu_{C}(A\cup B)\leq\mu_{C}(A)+\mu_{C}(B)\text{ with equality for }A\perp_{C}B,$ (subadditivity)
$\mu_{C}(A)\leq\mu_{D}(A)\text{ if }C<D,$ (insight monotonicity)
$\sup\{\mu_{C}(A)\mid A\subseteq\mathcal{K}\}<\infty,$ (boundedness)
@@
where $A\perp_{C}B$ means that no element of $A$ can derive elements of $B$ within computational budget $C$, and vice versa.
The idea of computational independence allows for distinguishing between facts derivable from other facts and facts that cannot be derived that way. We need that distinction to deal with the mechanics of reasoning process. Also, as we will see in the following sections, it plays an important role in our study of the semantic information conservation and properties of bounded reasoning.
Definition 4 (Computational Independence).
Let $\mathcal{F}_{C}$ denote the set of all computational transformations with complexity bounded by $C$. Knowledge sets $A,B\subseteq\mathcal{K}$ are computationally independent with respect to computational budget $C$, denoted $A\perp_{C}B$, if and only if for all $f\in\mathcal{F}_{C}$:
$f(A)\cap B=\emptyset\quad\text{and}\quad A\cap f(B)=\emptyset.$ (1)
We limit our investigations to those cases for which the following model applies. Knowledge that cannot resolve any uncertainty or knowledge irrelevant to a query ($A\cap q=\emptyset$) has zero information. If that is not the case and there is information within our reach that reduces uncertainty, we may hope for some content useful for reasoning. To make that possible, we need to allow for information ordering. First, we assume information content is additive for independent knowledge sets, i.e., when no knowledge in $A$ can be derived from $B$ within computational budget $C$, and vice versa. Second, with redundant information in $A\cup B$ , we admit potential loss of information that may be caused by aggregation. Finally, we say that with increasing computational budget we can extract more knowledge (ordered information) from the same information subset.
5.1.1 Justification for the Axioms of the Semantic Information Measure
The choice of conditional additivity for computationally independent sets, i.e., demanding $\mu_{C}(A\cup B)\leq\mu_{C}(A)+\mu_{C}(B)$ with equality for $A\perp_{C}B$, is necessary to describe LLM inference as an auction of ideas. As we will see, we need to isolate knowledge encoded in distinct neural circuits, attention heads, or activation patterns. The concept of Computational Independence is the formal translation of this informational separation into a computational context. It formally states that the knowledge in idea $A$ cannot be used to derive the knowledge in idea $B$ within a given computational budget $C$, and vice-versa.
Given this premise of informational isolation, the additivity of the semantic measure becomes the baseline definition of non-interaction. If two knowledge sets are truly independent, they offer no redundant or synergistic information. That baseline is necessary to identify and quantify deviations. As we will see in the following section, the aggregation of diverse probabilistic beliefs violates this additivity, representing a form of forced synergy imposed by the inference process.
This axiomatic choice is most applicable to the idealized setting of Theorem 6. Later, in Theorem 8 and 9, we relax that model by moving the source of interaction from the measure itself to the aggregation function in actual transformers.
5.1.2 Justification for the Axiom of Unlimited Reasoning with Emergence Operator
The completeness (or closure property) of the set of reasoning functions $\mathcal{F}_{\infty}$ sounds somewhat mysterious, but is rather natural. We demand that for any reasoning steps $f$ and $g$, there exists a single step $h$ such that $g(A\cup f(A))=h(A)$. That is the nature of an idealized, complete reasoning system we want to investigate to understand its tendency to generate hallucinations.
The construct $\mathcal{F}_{\infty}$ is designed to represent the theoretical limit of reasoning with an unbounded computational capacity. It is not intended to model the practical, heuristic, and bounded reasoning of a real LLM, but rather the theoretical limit of what is ultimately knowable from a set of premises. Completeness is fundamental to formal systems, such as first-order logic or Turing-complete computations and Direct Revelation Principle in auction theory kuratowski_set_1978 ; sipser1996introduction ; fudenberg1991game . If one can prove a lemma $f(A)$ from a set of axioms $A$, and then prove a theorem $g(A\cup f(A))$ using that lemma, a direct proof of the theorem from the axioms alone $h(A)$ must exist by simply concatenating the proof steps. The composition of two computable functions is itself a computable function. A direct revelation mechanism in (auction) game theory is a composition of preference processing mechanisms where agents report their preferences directly to the mechanism, and the mechanism implements their preference processing strategies to generate outcome.
That axiom is essential for the proof of idempotence of unlimited reasoning in Theorem 3 Theorem 4. These are critical philosophical and mathematical results necessary for our study. With them we argue that the information gain observed in practical, budget-constrained inference is an artifact of computational bounds, not the magical creation of knowledge ex nihilo.
5.1.3 Limits of Knowledge and Reasoning
If we accept that knowledge model, then we must also accept there exists a limit to its ability to resolve uncertainty.
Theorem 1 (Limits of Knowledge).
For any $A\subseteq\mathcal{P}(\mathcal{K})$ there exists
$\mu_{\infty}(A)=\lim_{C\to\infty}\mu_{C}(A).$ (2)
Proof.
Since $\mu_{C}(A)\leq\mu_{D}(A)$ when $C<D$ and $\mu_{C}$ is bounded on finite $\mathcal{K}$, the monotonic increasing and bounded sequence $\{\mu_{C}(A)\}_{C}$ converges with $C\to\infty$ for any subset of knowledge space. ∎
It may be surprising to discover that the existence of that limit does not prohibit emergent information. We can learn and obtain useful information without violating that fundamental limitation above. To explain that critical and controversial property of information processing, we first introduce the emergence operator $\mathcal{E}_{C}$ to describe knowledge becoming accessible through (computationally) bounded reasoning. Then we can study the implications in more details.